# The Rhaly Operators on Köthe Spaces

### v1

@@ 1. Introduction
is obtained. The Cesáro operator has been a subject of significant study within the classical sequence spaces $\ell_{p}$ ($1<p<\infty$), $\ell_{\infty}$, $c_{0}$, $c$, the weighted Banach spaces $\ell_{p}(w)$, $1<p<\infty$ and the weighted Banach spaces $c_{0}(w)$, $1<p<\infty$ for some references, we refer [1, 4, 5, 9, 12, 24].
Investigating the Cesáro operator on Fréchet spaces is also a topic of interest; we refer [6, 7]. In [7], Albanese, Bonet, and Ricker studied the Cesáro operator $\mathcal{C}$ on power series spaces of finite type $\Lambda_{1}(\alpha)$ and they focused on its spectrum, which significantly differs based on the nuclearity of $\Lambda_{1}(\alpha)$ and characterized the spectrum. Furthermore, they showed that $\mathcal{C}$ is power bounded, uniformly mean ergodic whenever $\Lambda_{1}(\alpha)$ is nuclear.
@@ 2. Preliminaries
  • power bounded if for every $p\in\mathbb{N}$ there exists $q\in\mathbb{N}$ such that
    $\|T^{k}(x)\|_{p}\leq\|x\|_{q}$
@@
  • Cesáro bounded if for every $p\in\mathbb{N}$ there exists $q\in\mathbb{N}$ such that
    $\|T^{[k]}(x)\|_{p}\leq\|x\|_{q}$
@@
    $T$ is power bounded if and only if for every $p\in\mathbb{N}$ there exists $q\in\mathbb{N}$ such that
    $\|T^{k}(e_{n})\|_{p}\leq\|e_{n}\|_{q}$
@@
    $T$ is Cesáro bounded if and only if for every $p\in\mathbb{N}$ there exists $q\in\mathbb{N}$ such that
    $\|T^{[k]}(e_{n})\|_{p}\leq\|e_{n}\|_{q}$
@@ 3. Rhaly Operators
$-k\beta_{n}\leq\frac{B}{A}-\frac{1}{mA}\alpha_{n}$
@@ 4. Integral Representation of Rhaly Operators
[…]
@@
[…]
@@ 5. Topologizability and Power Boundedness of Rhaly Operators
The topologizability and power boundedness of the Rhaly operator will be established for $\Lambda_{1}(\alpha)$, $\Lambda_{\infty}(\alpha)$, and for some Köthe spaces $K(a_{n,k})$. Consequently, these results will enable us to determine the properties of Rhaly operators on $H(\mathbb{C})$ and $H(\mathbb{D})$.
@@
Let $K(a_{n,k})$ be a Köthe space and let $\theta\in K(a_{n,k})$. Assume that $A=\inf_{n\in\mathbb{N}}a_{n,m_{0}}>0$ for some $m_{0}\in\mathbb{N}$ and $\sup_{p\in\mathbb{N}}\|\theta\|_{p}\leq A$. The Rhaly operator $R_{\theta}:K(a_{n,k})\to K(a_{n,k})$ is power bounded.
@@
Let $A=\inf_{n\in\mathbb{N}}a_{n,m_{0}}>0$ for some $m_{0}\in\mathbb{N}$ and $\theta\in K(a_{n,k})$ with $\sup_{p\in\mathbb{N}}\|\theta\|_{p}\leq A$. Then for every $p\in\mathbb{N}$, the numbers $D_{p}=\frac{\|\theta\|_{p}}{A}$ are less then or equal to 1. Inequality (5.1) says that $R_{\theta}$ is power bounded. ∎
@@
Let $K(a_{n,k})$ be a Montel Köthe space and let $\theta\in K(a_{n,k})$. Assume that $A=\inf_{n\in\mathbb{N}}a_{n,m_{0}}>0$ for some $m_{0}\in\mathbb{N}$ and $\sup_{p\in\mathbb{N}}\|\theta\|_{p}\leq A$. The Rhaly operator $R_{\theta}:K(a_{n,k})\to K(a_{n,k})$ is
@@
Let $\theta\in\Lambda_{\infty}(\alpha)$. Then for every $p\in\mathbb{N}$, there exists a $q_{p}\in\mathbb{N}$ satisfying $\|\theta\|_{p}\leq e^{q_{p}\alpha_{1}}.$ We proceed by following the steps outlined in the proof of Proposition 5.1. In this case, $m_{0}$ is choosen as $q_{1}$ and $A=\inf_{n\in\mathbb{N}}a_{n,q_{1}}=e^{q_{1}\alpha_{1}}>0$. We recall that $D_{p}=\frac{\|\theta\|_{p}}{A}$ for every $p\in\mathbb{N}$ and then we get $D_{1}\leq 1$. It can also be assumed that $e^{q_{1}\alpha_{1}}\geq 1$. By using (5.1), we obtain
$\begin{split}\|R^{k}_{\theta}e_{n}\|_{p}&\leq D_{p}(D_{1})^{k-1}\|e_{n}\|_{q_{1}}\leq\frac{\|\theta\|_{p}}{e^{q_{1}\alpha_{1}}}e^{q_{1}\alpha_{n}}\\ &\leq\|\theta\|_{p}e^{q_{1}\alpha_{n}}\leq e^{(q_{p}+q_{1})\alpha_{n}}=\|e_{n}\|_{q_{p}+q_{1}}\end{split}$
By Lemma 2.4, $R_{\theta}:\Lambda_{\infty}(\alpha)\to\Lambda_{\infty}(\alpha)$ is power bounded. ∎
@@
for every $m\geq n$, $k,n\in\mathbb{N}$. Hence, we obtain that
$|(R^{k}_{\theta}e_{n})_{m}|\leq\left(\sum^{m}_{j=n}|\theta_{j}|\right)^{k}\hskip 10.84006pt\text{and}\hskip 10.84006pt\|R^{k}_{\theta}e_{n}\|_{p}\leq\sup_{m\geq n}\left(\sum^{m}_{j=n}|\theta_{j}|^{k}\right)e^{-\frac{1}{p}\alpha_{m}}$
@@
$\begin{split}\|R^{k}_{\theta}e_{n}\|_{p}&\leq\sup_{m\geq n}\left(\sum^{m}_{j=n}|\theta_{j}|^{k}\right)e^{-\frac{1}{p}\alpha_{m}}\\ &=\sup_{m\geq n}\left(\sum^{m}_{j=n}|\theta_{j}|^{k}\right)e^{-\frac{1}{3p}\alpha_{m}}e^{-\frac{1}{3p}\alpha_{m}}e^{-\frac{1}{3p}\alpha_{m}}\\ &\leq e^{-\frac{1}{3p}\alpha_{n}}\sup_{m\geq n}\left(\sum^{m}_{j=n}|\theta_{j}|e^{-\frac{1}{3pk}\alpha_{j}}\right)^{k}e^{-\frac{1}{3p}\alpha_{m}}\\ &\leq e^{-\frac{1}{3p}\alpha_{n}}\left(\sum^{\infty}_{j=n}|\theta_{j}|e^{-\frac{1}{3pk}\alpha_{j}}\right)^{k}\sup_{m\geq n}e^{-\frac{1}{3p}\alpha_{m}}\\ &\leq e^{-\frac{1}{3p}\alpha_{n}}\left(\|\theta\|_{3pk}\right)^{k}=\|e_{n}\|_{3p}\left(\|\theta\|_{3pk}\right)^{k}\end{split}$ (5.2)
@@
Let $\Lambda_{1}(\alpha)$ be a nuclear space and $\theta\in\Lambda_{1}(\alpha)$. The Rhaly operator $R_{\theta}:\Lambda_{1}(\alpha)\to\Lambda_{1}(\alpha)$ is topologizable. If $\sup_{p\in\mathbb{N}}\|\theta\|_{p}<\infty$, $R_{\theta}$ is m-topologizable.
Theorem 5.9.
Let $\Lambda_{1}(\alpha)$ be a nuclear space and $\theta\in\Lambda_{1}(\alpha)$. The Rhaly operator $R_{\theta}:\Lambda_{1}(\alpha)\to\Lambda_{1}(\alpha)$ is power bounded if and only if $\sup_{p\in\mathbb{N}}\|\theta\|_{p}\leq 1$.
Proof.
Let assume that $\sup_{p\in\mathbb{N}}\|\theta\|_{p}\leq 1$. (5.2) gives us that for every $p,k,n\in\mathbb{N}$, it yields that
$\|R^{k}_{\theta}e_{n}\|_{p}\leq\|e_{n}\|_{3p}.$
Then $R_{\theta}$ is power bounded. For the converse, we assume that $R_{\theta}:\Lambda_{1}(\alpha)\to\Lambda_{1}(\alpha)$ is power bounded. This means that for every $p\in\mathbb{N}$ there exists a $q\in\mathbb{N}$ such that
$\|R^{k}_{\theta}e_{n}\|_{p}\leq\|e_{n}\|_{q}$
holds for every $k,n\in\mathbb{N}$. By choosing $n=k=1$, we have
$\|\theta\|_{p}=\|R_{\theta}e_{1}\|_{p}\leq\|e_{1}\|_{q}=e^{-\frac{1}{q}\alpha_{1}}<1.$
This completes the proof. ∎
@@
Let $\Lambda_{1}(\alpha)$ be a nuclear power series space of finite type. The following properties are equivalent for the Rhaly operator $R_{\theta}:\Lambda_{1}(\alpha)\to\Lambda_{1}(\alpha)$, where $\theta\in\Lambda_{1}(\alpha)$;
@@
  • 5.
    $\sup_{p\in\mathbb{N}}\|\theta\|_{p}\leq 1$
Proof.
Theorem 5.9 asserts the equivalence between conditions (1) and (5). Theorem 2.6 (by Albanese, Bonet, and Ricker) regarding the mean ergodicity of Montel Köthe spaces, establishes that (1) implies (2). Furthermore, Theorem 2.7 (by Kalmes and Santacreu) tells us that statements (2), (3), and (4) are equivalent. We now turn to proving that (4) implies (5). Let us assume that $R_{\theta}$ is Cesáro bounded. Then, for every $p\in\mathbb{N}$, there exists a $q\in\mathbb{N}$ such that for every $k,n\in\mathbb{N}$
$\|R^{[k]}_{\theta}(e_{n})\|_{p}\leq\|e_{n}\|_{q}$
holds. In particular, taking $k=n=1$, we get
$\|\theta\|_{p}=\|R_{\theta}(e_{1})\|_{p}\leq\|e_{1}\|_{q}=e^{-\frac{1}{q}\alpha_{1}}<1$
which shows that (5) is satisfied. This completes the proof. ∎
@@
The following statements are equivalent for the operator $R_{\theta}:H(\mathbb{D})\to H(\mathbb{D})$:

### v2

@@ 1. Introduction
is obtained. The Cesáro operator has been a subject of significant study within the classical sequence spaces $\ell_{p}$ ($1<p<\infty$), $\ell_{\infty}$, $_{0}, cc$, the weighted Banach spaces $\ell_{p}(w)$, $1<p<\infty$ and the weighted Banach spaces $c_{0}(w)$, $1<p<\infty$ for some references, we refer [1, 4, 5, 9, 12, 18, 28].
Investigating the Cesáro operator on Fréchet spaces is also a topic of interest; we refer [6, 7, 20, 21, 22]. In [7], Albanese, Bonet, and Ricker studied the Cesáro operator $\mathcal{C}$ on power series spaces of finite type $\Lambda_{1}(\alpha)$ and they focused on its spectrum, which significantly differs based on the nuclearity of $\Lambda_{1}(\alpha)$ and characterized the spectrum. Furthermore, they showed that $\mathcal{C}$ is power bounded, uniformly mean ergodic whenever $\Lambda_{1}(\alpha)$ is nuclear.
@@ 2. Preliminaries
  • power bounded if for every $p\in\mathbb{N}$ there exist $q\in\mathbb{N}$ and $C_{p}\geq 1$ such that
    $\|T^{k}(x)\|_{p}\leq C_{p}\|x\|_{q}$
@@
  • Cesáro bounded if for every $p\in\mathbb{N}$ there exist $q\in\mathbb{N}$ and $C_{p}\geq 1$ such that
    $\|T^{[k]}(x)\|_{p}\leq C_{p}\|x\|_{q}$
@@
    $T$ is power bounded if and only if for every $p\in\mathbb{N}$ there exist $q\in\mathbb{N}$ and $C_{p}\geq 1$ such that
    $\|T^{k}(e_{n})\|_{p}\leq C_{p}\|e_{n}\|_{q}$
@@
    $T$ is Cesáro bounded if and only if for every $p\in\mathbb{N}$ there exist $q\in\mathbb{N}$ and $C_{p}$ such that
    $\|T^{[k]}(e_{n})\|_{p}\leq C_{p}\|e_{n}\|_{q}$
@@ 3. Rhaly Operators
$k\beta_{n}\leq\frac{B}{A}-\frac{1}{mA}\alpha_{n}$
@@ 4. Integral Representation of Rhaly Operators
$(R_{g}f)(z)=\frac{1}{2\pi i}\int_{|w|=r_{0}}\frac{f(w)}{w(1-w)}g\left(\frac{z}{w}\right)dw=\sum^{\infty}_{n=0}\left(\sum^{n}_{k=0}a_{k}\right)b_{n}z^{n}$
@@
$\begin{split}&\frac{1}{2\pi i}\int_{|w|=r_{0}}\frac{f(w)}{w(1-w)}g\left(\frac{z}{w}\right)dw=\frac{1}{2\pi i}\int_{|w|=r_{0}}\frac{h(w)}{w}g\left(\frac{z}{w}\right)dw\\ &=\sum^{\infty}_{n=0}b_{n}z^{n}\frac{1}{2\pi i}\int_{|w|=r_{0}}\frac{h(w)}{w^{n+1}}dw=\sum^{\infty}_{n=0}c_{n}b_{n}z^{n}=\sum^{\infty}_{n=0}\left(\sum^{n}_{k=0}a_{k}\right)b_{n}z^{n}\end{split}$
@@ 5. Topologizability and Power Boundedness of Rhaly Operators
The topologizability and power boundedness of the Rhaly operator will be established for $\Lambda_{1}(\alpha)$, $\Lambda_{\infty}(\alpha)$, and for some Köthe spaces $K(a_{n,k})$. Consequently, these results will enable us to determine the properties of Rhaly operators on the spaces $H(\mathbb{C})$ and $H(\mathbb{D})$.
@@
Let $K(a_{n,k})$ be a Köthe space and let $\theta\in K(a_{n,k})$. Assume that $A=\inf_{n\in\mathbb{N}}a_{n,m_{0}}>0$ for some $m_{0}\in\mathbb{N}$ and $\|\theta\|_{m_{0}}\leq A$. The Rhaly operator $R_{\theta}:K(a_{n,k})\to K(a_{n,k})$ is power bounded.
@@
Let $A=\inf_{n\in\mathbb{N}}a_{n,m_{0}}>0$ for some $m_{0}\in\mathbb{N}$ and $\theta\in K(a_{n,k})$ with $\|\theta\|_{m_{0}}\leq A$. Then the number $D_{m_{0}}=\frac{\|\theta\|_{m_{0}}}{A}$ is less then or equal to 1. Inequality (5.1) says that $R_{\theta}$ is power bounded. ∎
@@
Let $K(a_{n,k})$ be a Montel Köthe space and let $\theta\in K(a_{n,k})$. Assume that $A=\inf_{n\in\mathbb{N}}a_{n,m_{0}}>0$ for some $m_{0}\in\mathbb{N}$ and $\|\theta\|_{m_{0}}\leq A$. The Rhaly operator $R_{\theta}:K(a_{n,k})\to K(a_{n,k})$ is
@@
Let $\theta\in\Lambda_{\infty}(\alpha)$. Then for every $p\in\mathbb{N}$, there exists a $q_{p}\in\mathbb{N}$ satisfying $\|\theta\|_{p}\leq e^{q_{p}\alpha_{1}}=\inf_{n\in\mathbb{N}}e^{q_{p}\alpha_{n}}.$ By Proposition 5.3, we obtain that $R_{\theta}:\Lambda_{\infty}(\alpha)\to\Lambda_{\infty}(\alpha)$ is power bounded. ∎
@@
for every $m\geq n$, and 0 for every $m<n$, $m,k,n\in\mathbb{N}$ . Hence, we obtain that
$|(R^{k}_{\theta}e_{n})_{m}|\leq\left(\sum^{m}_{j=n}|\theta_{j}|\right)^{k}\hskip 10.84006pt\text{and}\hskip 10.84006pt\|R^{k}_{\theta}e_{n}\|_{p}\leq\sup_{m\geq n}\left(\sum^{m}_{j=n}|\theta_{j}|\right)^{k}e^{-\frac{1}{p}\alpha_{m}}$
@@
$\begin{split}\|R^{k}_{\theta}e_{n}\|_{p}&\leq\sup_{m\geq n}\left(\sum^{m}_{j=n}|\theta_{j}|\right)^{k}e^{-\frac{1}{p}\alpha_{m}}\\ &=\sup_{m\geq n}\left(\sum^{m}_{j=n}|\theta_{j}|\right)^{k}e^{-\frac{1}{3p}\alpha_{m}}e^{-\frac{1}{3p}\alpha_{m}}e^{-\frac{1}{3p}\alpha_{m}}\\ &\leq e^{-\frac{1}{3p}\alpha_{n}}\sup_{m\geq n}\left(\sum^{m}_{j=n}|\theta_{j}|e^{-\frac{1}{3pk}\alpha_{j}}\right)^{k}e^{-\frac{1}{3p}\alpha_{m}}\\ &\leq e^{-\frac{1}{3p}\alpha_{n}}\left(\sum^{\infty}_{j=n}|\theta_{j}|e^{-\frac{1}{3pk}\alpha_{j}}\right)^{k}\sup_{m\geq n}e^{-\frac{1}{3p}\alpha_{m}}\\ &\leq e^{-\frac{1}{3p}\alpha_{n}}\left(\|\theta\|_{3pk}\right)^{k}=\|e_{n}\|_{3p}\left(\|\theta\|_{3pk}\right)^{k}\end{split}$ (5.2)
@@
Let $\Lambda_{1}(\alpha)$ be a nuclear space and $\theta\in\Lambda_{1}(\alpha)$.
  • i.
    The Rhaly operator $R_{\theta}:\Lambda_{1}(\alpha)\to\Lambda_{1}(\alpha)$ is topologizable.
  • ii.
    If $\sup_{p\in\mathbb{N}}\|\theta\|_{p}=\sup_{n\in\mathbb{N}}|\theta_{n}|<\infty$, that is, $\theta\in\ell_{\infty}$, then $R_{\theta}$ is m-topologizable.
  • iii.
    If $\sup_{p\in\mathbb{N}}\|\theta\|_{p}=\sup_{n\in\mathbb{N}}|\theta_{n}|\leq 1$, then $R_{\theta}$ is power bounded.
@@
Let $\Lambda_{1}(\alpha)$ be a nuclear power series space of finite type and $\theta\in\Lambda_{1}(\alpha)$.Suppose that $\sup_{p\in\mathbb{N}}\|\theta\|_{p}=\sup_{n\in\mathbb{N}}|\theta|_{n}\leq 1$. Then the Rhaly operator $R_{\theta}:\Lambda_{1}(\alpha)\to\Lambda_{1}(\alpha)$ satisfies the following:
@@
The Rhaly operator $R_{g}:H(\mathbb{D})\to H(\mathbb{D})$ is